# Proposal for DESC cohesive Deep Drilling Programs for LSST A Cohesive Deep Drilling Field Strategy for LSST Cosmology

### v1

@@ 1 Introduction
The Vera C. Rubin Observatory Legacy Survey of Space and Time (LSST; Ivezić et al. (2019)) will image billions of objects in six bands for ten years. The observing time will be shared between the wide-fast-deep (WFD) primary survey, which will cover half of the sky (20,000 deg2), and a set of minisurveys, including intensive observation of a set of deep drilling fields (DDFs). In the WFD survey, fields are observed with a similar cadence and pattern, the cadence beeing defined as the median internight gap in any filter (high cadences correspond to small internight gaps). The DDF mini-survey is characterized by a high cadence of observation, a high number of visits per observing night, and a limited survey area (60-70 deg2).
The DDF mini-survey is vital for achieving the core science the LSST Dark Energy Science Collaboration (DESC; Abate (2012)) and it has three main drivers: cosmology, calibration of primary systematic uncertainties for several cosmological measurements, and synergy with other surveys. DDFs are essential for DESC to reach its Stage IV dark energy goals, that is to multiply by 10 the Figure of Merit (FoM) of stage II surveys (Albrecht et al., 2006).
DESC has long emphasized the importance of the observing strategy for science probes (or key input to science probes) used to measure cosmological parameters (dark energy equation of state) with high accuracy. Many studies have already been achieved (Awan et al. (2016), Lochner et al. (2018), Scolnic et al. (2018), Almoubayyed et al. (2020), Lochner et al. (2021a), Lochner et al. (2021b), Gris et al. (2023), Alves et al. (2023)) and the work presented in this paper is a continuation of many years of efforts aiming to define the optimal observing strategy to accomplish the scientific objectives of the Rubin Observatory. In the following we focus on the design of the DDF mini-survey which is critical for three DESC cosmological measurements: photometric redshifts, weak gravitational lensing, and Type Ia Supernovae, as outlined in the text that follows.
@@
Weak gravitational lensing (WL), the deflection of light from distant massive structures, is currently one of the most promising probes to constrain the growth of the cosmic structure and unveil the nature of dark energy (Mandelbaum (2018) and references therein). The LSST survey will dramatically increase the statistical power of WL measurements by observing billions of galaxies. Two key parameters are to be extracted from LSST data: the shear from galaxy shapes, and the photometric redshift estimates. Shear signals are to be measured with systematic errors lower than to minimize the degradation in cosmological parameter accuracies (Huterer et al., 2006). The shear calibration bias, one of the largest systematic error sources, may be reduced using self-calibration techniques such as metacalibration (Huff & Mandelbaum (2017), Sheldon & Huff (2017)) where real images are used to simulate the effect of a known shear. Using the metacalibration method in images collected in the WFD survey would lead to a loss of precision in the statistical shear measurements (Zhang et al., 2023). Reducing this loss (pixel noise) to 5 in metacalibration estimators of weak lensing signals requires using the DDFs and the deep-field metacalibration technique (Zhang et al., 2023).
@@
Type Ia Supernovae (SNe Ia) are the results of powerful and luminous explosions of white dwarfs. SNe Ia can be considered as standardizable candles to estimate distances used to measure cosmological parameters (equation of state of dark energy) in a Hubble diagram (Betoule et al. (2014), Brout et al. (2022)). Cosmological distances are estimated from SNe Ia parameters extracted from photometric observations (SNe Ia light curves) generated by the observing strategy. Hence measuring cosmological parameters with a high accuracy requires optimizing the observing strategy to collect high quality light curves (in terms of sampling and signal-to-noise ratio). LSST will observe a large number of SNe Ia at low redshifts (0.4) in the WFD survey (Lochner et al., 2021a). Measuring and parameters of the equation of state of dark energy requires to observe (necessary condition) a large sample of SNe Ia at higher redshifts (up to 1.1) and this is only feasible with DDFs (Gris et al., 2023). DDFs are thus critical if LSST is expected to address dark energy physics and cosmology with SNe Ia.
@@
The purpose of this paper is to design a DESC cohesive DDF program that would fulfill calibration requirements (photo- and WL) and cosmological constraints (SNe Ia) whilst meeting the Survey Cadence Optimization Committee (SCOC; Bianco et al. (2021)) phase 2 guidelines (The Rubin Observatory Survey Cadence Optimization Committee, 2022). This article is divided in four sections. The design requirements of the survey are detailed in §2 with a summary of the SCOC phase 2 recommendations (§2.1) and a detailed explanation of the requirements from cosmology measurements (§2.2). A set of metrics to assess observing strategy and to quantify to which extent science requirements are met are defined in §3. These metrics are based on statistical optimization and do not include systematic uncertainties. A proposed cohesive DDF strategy is presented in §4 along with the methodology used to define observing strategy parameters leading to DDF surveys compliant with science requirements. These strategies are simulated and assessed in §5 using the metrics defined in §3.
@@ 2.2.1 Photo-
It may be possible to break degeneracies by combining wide-field, multi-band measurements with accurate redshift galaxies. In deep-fields, precise photometry of multi-band flux can be used to group galaxies into fine-grained set of phenotypes (Sánchez & Bernstein, 2018), which are types of galaxy based on observed flux rather than on rest-frame properties. The multi-band deep fields provide greater opportunities to assess the impact of selection effects.The set of galaxy photometry and a set of galaxies with confidently known (obtained using many-band photometry observations or spectroscopy) can be used as input to a photometric method such as SOMPZ which is based upon the self-organizing map algorithm (Buchs et al. (2019), Myles et al. (2021)).
@@ 3.1 Photo-
Photo- calibration requirements can be translated to 5- depth point source constraints (subsubsection 2.2.1). We define the photo- metric as the (coadded per band) 5- depth difference between observing strategy observations and the requirements defined in Table 2:
@@ 3.2 Weak Lensing
Following the WL requirements (subsubsection 2.2.2) we define the photo- metric as:
where is the number of visits (per band) of the proposed strategies and the number of visits corresponding to WL requirements (Table 2). WL requirements are fulfilled if 1.
@@ 3.3 SNe Ia
where , where is the overall flux normalization, describes the width of the light curve, is equal to a color offset (with respect to the average) at the date of peak brightness in -band, . For each SN Ia, parameters are estimated from a fit of a SN Ia model to the measurements of a multicolor light curve. , and are global parameters estimated from the data. is the absolute magnitude in rest-frame -band for a median SN Ia with (,) = (0.0, 0.0). and are global nuisance parameters quantifying the correlation between brightness with and , respectively. The three parameters are usually fitted along with cosmological parameters by minimizing Equation 10. In this study we set , and to their simulated values so as to fit only cosmological parameters.
In the SALT3 model a SNe Ia is described by five parameters: , , , (redshift), and (time of maximum luminosity). For each of the proposed strategies we have simulated a sample of SNe Ia corresponding to 30 times the number of expected SNe Ia (we use the rate Hounsell et al. (2018)) in the full redshift range [0.01,1.1]. Concerning the (,) distribution of SNe Ia, we use the G10 intrinsic scatter model (Scolnic & Kessler, 2016) where (,) distributions are described by asymmetric gaussian distributions with three parameters. We choose random values spanning over the season duration of a group of observations.
@@
One may notice that the SMoM metric is estimated by combining SNe Ia samples observed in the WFD (low-) and DDF (high-) surveys. It can not be used to check directly whether SNe Ia specifications (subsubsection 2.2.3) related to DDFs are fulfilled. But it has been shown that UDF scenarios with corresponding to SNe Ia requirements tend to maximize SMoM values (Gris et al., 2023).
@@ 4 Designing a cohesive DDF strategy
As shown in §2, measuring cosmological parameters (dark energy equation of state) with high accuracy with SNe Ia requires to observe (necessary condition) a large sample of SNe Ia at higher redshifts (0.8). This can be achieved with a high cadence of observation ( 2 night) and a large number of visits per observing night (more than 150 visits for a 2 night cadence) with a specific filter allocation defined by the -depth to be reached (see Gris et al. (2023) for all the details). Nonetheless, this corresponds to a total number of visits (per season) quite costly in terms of budget: observing a SNe Ia sample up to a redshift completeness of 0.75 leads to a fraction of DD visits of about 0.9% per season (Table 3). Since the DDF budget is limited to 5-7%, it is not realistic to design a uniform DD survey (leading to accurate cosmological measurements) where all the fields would be observed at the same cadence, with the same number of visits per observing night, and the same filter allocation, for ten years.
@@
It is possible from Equation 15 to estimate the parameter space (N, N) for a choice of (N,N) configurations if all other parameters are assigned the values of Table 5. Since SNe Ia requirements are related to the number of visits per observing night (subsubsection 2.2.3), we define a DDF strategy by the four parameters (N/obs. night, N, N, N). This requires to choose a cadence of observation (cad) and season length (sl) for UD fields. We have used cad=2 nights and sl=180 days in this paper.
@@
Completing the definition of the DESC cohesive DDF strategies requires to choose three additional parameters: season length, cadence of observation, and filter allocation (number of visits per band and per observing night). For UDFs, these parameters are coming from depth requirements related to the redshift completeness of the SNe Ia survey (§2.2 and Table 5). For DFs, we have chosen a season length of 180 days (to maximize the number of observed SNe Ia ) and a cadence of three days (median cadence observed in the simulations). The number of visits per band and per observing night (filter allocation) for the DFs is estimated from the total number of visits per season N/season (Figure 1). N/season is distributed among filters and observing nights in such a way that photo- and WL requirements are fulfilled (subsubsection 2.2.1 and subsubsection 2.2.2). If this is not possible (which is the case for all the proposed scenarios except the ones optimal for photo-/WL) the number of visits is split in such a way that the relative number of visits per band is equal to the one estimated from photo- and WL requirements.
The LSST camera has 6 filters () and only five can be loaded in the filter exchange carousel. The current filter load strategy is to replace during daytime one of the or by the band at the start of dark time, when the lunar phase is lower than 20%. The process is reversed when the lunar phase is above 20%. We have performed a study using SNe Ia to assess the impact of the swap or on two metrics ( and N). The results are detailed in Appendix A . The conclusion is that the minimal loss is observed for the swap. This can be explained by the fact that the number of visits in the band is lower than the number of visits in the band. The decrease of the light curve signal-to-noise ratio due to the swap is lower if and bands are exchanged. Following the conclusion of this study we have swapped the and bands when the lunar phase is low in the design of the proposed strategies. The total number of () visits is distributed among the nights with a Moon phase lower (higher) than 20%.
@@ 5.3 WL metric
Season 1 results (Tab. 13) show that WL requirements are met for all fields except for ECDFS-a and ECDFS-b in the southern region. The number of visits in ,,, for season 1 is driven by WL requirements (Tab. 2). is thus close to one for these bands and for all fields but EDFS-a and EDFS-b. For these two fields the number of visits is a factor of 2 lower compared to the other DDFs (SCOC phase 2 recommendations) and the ratio is close to 0.5. For this season the requirement 16 (subsubsection 2.2.2) is fullfilled for all the fields but EDFS-a and EDFS-b and only for and bands.
@@ 5.4 SNe Ia metric
We assume that the SNe Ia used to estimate SMoM have host galaxy redshifts measured from spectroscopic observations. We (conservatively) consider two spectroscopic resources contemporaneous with LSST to provide vital spectra and host redshifts, PFS/Subaru (Tamura et al., 2016) and 4MOST/TiDES (de Jong et al. (2019); Swann et al. (2019)). The PFS/Subaru spectroscopic follow-up survey is designed to observe two of the northernmost LSST DDFs, COSMOS and XMM-LSS. About 20,000 host galaxy redshifts up to 0.8 will be measured after ten years. The 4MOST/TiDES survey is dedicated to the spectroscopic follow-up of extragalactic optical transients. About 50,000 host galaxy redshifts up to 1 will be collected for 5 years. This corresponds both to the WFD and the DD fields. The fraction of SNe Ia expected to have secure redshift measurements is taken from Mandelbaum et al. (2019) (Figure 1, right plot).
@@
Measuring cosmological parameters () with high accuracy (i.e. maximizing SMoM) requires to observe (necessary condition) a large sample of well-measured SNe Ia at higher redshifts (Gris et al., 2023). Scenarios with UDFs lead (by design) to samples with 1000, 1500, 1800 SNe Ia with for DDF_DESC_0.80*, DDF_DESC_0.75*, DDF_DESC_0.70* observing strategies, respectively (Figure 12). Furthermore, a large fraction of SNe Ia in these samples (65-80, 40-60, 20-40, respectively) lead to accurate distance measurements ().

### v2

@@ 1 Introduction
The Vera C. Rubin Observatory Legacy Survey of Space and Time (LSST; Ivezić et al., 2019) will image billions of objects in six bands for ten years. The observing time will be shared between the wide-fast-deep (WFD) primary survey, which will cover half of the sky (20,000 deg2), and a set of minisurveys, including intensive observation of a set of deep drilling fields (DDFs). In the WFD survey, fields are observed with a similar cadence and pattern, the cadence beeing defined as the median internight gap in any filter (high cadences correspond to small internight gaps). The DDF mini-survey is characterized by a high cadence of observation, a high number of visits per observing night, and a limited survey area (60-70 deg2).
The DDF mini-survey is vital for achieving the core science the LSST Dark Energy Science Collaboration (DESC; Abate, 2012) and it has three main drivers: cosmology, calibration of primary systematic uncertainties for several cosmological measurements, and synergy with other surveys. DDFs are essential for DESC to reach its Stage IV dark energy goals, that is to multiply by 10 the Figure of Merit (FoM) of stage II surveys (Albrecht et al., 2006).
DESC has long emphasized the importance of the observing strategy for science probes (or key input to science probes) used to measure cosmological parameters (dark energy equation of state) with high accuracy. Many studies have already been achieved (Awan et al., 2016; Lochner et al., 2018; Scolnic et al., 2018; Almoubayyed et al., 2020; Lochner et al., 2021a, b; Gris et al., 2023; Alves et al., 2023) and the work presented in this paper is a continuation of many years of efforts aiming to define the optimal observing strategy to accomplish the scientific objectives of the Rubin Observatory. In the following we focus on the design of the DDF mini-survey which is critical for three DESC cosmological measurements: photometric redshifts, weak gravitational lensing, and Type Ia Supernovae, as outlined in the text that follows.
@@
Weak gravitational lensing (WL), the deflection of light from distant massive structures, is currently one of the most promising probes to constrain the growth of the cosmic structure and unveil the nature of dark energy (Mandelbaum, 2018, and references therein). The LSST survey will dramatically increase the statistical power of WL measurements by observing billions of galaxies. Two key parameters are to be extracted from LSST data: the shear from galaxy shapes, and the photometric redshift estimates. Shear signals are to be measured with systematic errors lower than to minimize the degradation in cosmological parameter accuracies (Huterer et al., 2006). The shear calibration bias, one of the largest systematic error sources, may be reduced using self-calibration techniques such as metacalibration (Huff & Mandelbaum, 2017; Sheldon & Huff, 2017) where real images are used to simulate the effect of a known shear. Using the metacalibration method in images collected in the WFD survey would lead to a loss of precision in the statistical shear measurements (Zhang et al., 2023). Reducing this loss (pixel noise) to 5 in metacalibration estimators of weak lensing signals requires using the DDFs and the deep-field metacalibration technique (Zhang et al., 2023).
@@
Type Ia Supernovae (SNe Ia) are the results of powerful and luminous explosions of white dwarfs. SNe Ia can be considered as standardizable candles to estimate distances used to measure cosmological parameters (equation of state of dark energy) in a Hubble diagram (Betoule et al., 2014; Brout et al., 2022). Cosmological distances are estimated from SNe Ia parameters extracted from photometric observations (SNe Ia light curves) generated by the observing strategy. Hence measuring cosmological parameters with a high accuracy requires optimizing the observing strategy to collect high quality light curves (in terms of sampling and signal-to-noise ratio). LSST will observe a large number of SNe Ia at low redshifts (0.4) in the WFD survey (Lochner et al., 2021a). Measuring and parameters of the equation of state of dark energy requires to observe a large sample of SNe Ia at higher redshifts (up to 1.1). This is a necessary but not sufficient condition to measure (,) with high accuracy because low-z SNe Ia are also required and this is only feasible with DDFs (Gris et al., 2023). DDFs are thus critical if LSST is expected to address dark energy physics and cosmology with SNe Ia.
@@
The purpose of this paper is to design a DESC cohesive DDF program that would fulfill calibration requirements (photo- and WL) and cosmological constraints (SNe Ia) whilst meeting the Survey Cadence Optimization Committee (SCOC; Bianco et al., 2021) phase 2 guidelines (The Rubin Observatory Survey Cadence Optimization Committee, 2022). This article is divided in four sections. The design requirements of the survey are detailed in §2 with a summary of the SCOC phase 2 recommendations (§2.1) and a detailed explanation of the requirements from cosmology measurements (§2.2). A set of metrics to assess observing strategy and to quantify to which extent science requirements are met are defined in §3. These metrics are based on statistical optimization and do not include systematic uncertainties. A proposed cohesive DDF strategy is presented in §4 along with the methodology used to define observing strategy parameters leading to DDF surveys compliant with science requirements. These strategies are simulated and assessed in §5 using the metrics defined in §3.
@@ 2.2.1 Photo-
It may be possible to break degeneracies by combining wide-field, multi-band measurements with accurate redshift galaxies. In deep-fields, precise photometry of multi-band flux can be used to group galaxies into fine-grained set of phenotypes (Sánchez & Bernstein, 2018), which are types of galaxy based on observed flux rather than on rest-frame properties. The multi-band deep fields provide greater opportunities to assess the impact of selection effects.The set of galaxy photometry and a set of galaxies with confidently known (obtained using many-band photometry observations or spectroscopy) can be used as input to a photometric method such as SOMPZ which is based upon the self-organizing map algorithm (Buchs et al., 2019; Myles et al., 2021).
@@ 3.1 Photo-
Photo- calibration requirements can be translated to 5- depth point source constraints (§2.2.1). We define the photo- metric as the (coadded per band) 5- depth difference between observing strategy observations and the requirements defined in Table 2:
@@ 3.3 SNe Ia
where , is the overall flux normalization, describes the width of the light curve, and is equal to a color offset (with respect to the average) at the date of peak brightness in -band, . For each SN Ia, parameters are estimated from a fit of a SN Ia model to the measurements of a multicolor light curve. , and are global parameters estimated from the data. is the absolute magnitude in rest-frame -band for a median SN Ia with (,) = (0.0, 0.0). and are global nuisance parameters quantifying the correlation between brightness with and , respectively. The three parameters are usually fitted along with cosmological parameters by minimizing Equation 10. In this study we set , and to their simulated values so as to fit only cosmological parameters.
In the SALT3 model a SNe Ia is described by five parameters: , , , (redshift), and (time of maximum luminosity). For each of the proposed strategies we have simulated a sample of SNe Ia corresponding to 30 times the number of expected SNe Ia (we use the SNe Ia production rate estimated in Hounsell et al., 2018) in the full redshift range [0.01,1.1]. Concerning the (,) distribution of SNe Ia, we use the G10 intrinsic scatter model (Scolnic & Kessler, 2016) where (,) distributions are described by asymmetric gaussian distributions with three parameters. We choose random values spanning over the season duration of a group of observations.
@@
One may notice that the SMoM metric is estimated by combining SNe Ia samples observed in the WFD (low-) and DDF (high-) surveys. It can not be used to check directly whether SNe Ia specifications (§2.2.3) related to DDFs are fulfilled. But it has been shown that UDF scenarios with corresponding to SNe Ia requirements tend to maximize SMoM values (Gris et al., 2023).
The SMoM metric does not include additional scatter and systematics effects (due to the degeneracy between the nuisance parameters and the intrinsic color of the SNe Ia distribution) that are to be taken into account if a measurement (Figure of Merit) is to be made. The main goal of the metric is give a first ranking of the proposed observing strategies using relative SMoM values. Estimating the Figure of Merit values for these strategies will require to include systematic uncertainties, to use realistic observing strategies (i.e. including realistic observing conditions) and to fit the nuisance parameters along with cosmological parameters. We do note, however, that adding systematic uncertainties would weaken any constraints, and so this optimization is a conservative approach to ensure that we aren’t dominated by the observing strategy systematics.
@@ 4 Designing a cohesive DDF strategy
As shown in §2, measuring cosmological parameters (dark energy equation of state) with high accuracy with SNe Ia requires to observe a large sample of SNe Ia at higher redshifts (0.8). This is a necessary but not sufficient condition to measure (,) with high accuracy because low-z SNe Ia are also required.This can be achieved with a high cadence of observation ( 2 night) and a large number of visits per observing night (more than 150 visits for a 2 night cadence) with a specific filter allocation defined by the -depth to be reached (see Gris et al., 2023, for all the details). Nonetheless, this corresponds to a total number of visits (per season) quite costly in terms of budget: observing a SNe Ia sample up to a redshift completeness of 0.75 leads to a fraction of DD visits of about 0.9% per season (Table 3). Since the DDF budget is limited to 5-7%, it is not realistic to design a uniform DD survey (leading to accurate cosmological measurements) where all the fields would be observed at the same cadence, with the same number of visits per observing night, and the same filter allocation, for ten years.
@@
It is possible from Equation 15 to estimate the parameter space (N, N) for a choice of (N,N) configurations if all other parameters are assigned the values of Table 5. Since SNe Ia requirements are related to the number of visits per observing night (§2.2.3), we define a DDF strategy by the four parameters (N/obs. night, N, N, N). This requires to choose a cadence of observation (cad) and season length (sl) for UD fields. We have used cad=2 nights and sl=180 days in this paper.
@@
Completing the definition of the DESC cohesive DDF strategies requires to choose three additional parameters: season length, cadence of observation, and filter allocation (number of visits per band and per observing night). For UDFs, these parameters are coming from depth requirements related to the redshift completeness of the SNe Ia survey (§2.2 and Table 5). For DFs, we have chosen a season length of 180 days (to maximize the number of observed SNe Ia ) and a cadence of three days (median cadence observed in the simulations). The number of visits per band and per observing night (filter allocation) for the DFs is estimated from the total number of visits per season N/season (Figure 1). N/season is distributed among filters and observing nights in such a way that photo- and WL requirements are fulfilled (§2.2.1 and §2.2.2). If this is not possible (which is the case for all the proposed scenarios except the ones optimal for photo-/WL) the number of visits is split in such a way that the relative number of visits per band is equal to the one estimated from photo- and WL requirements.
The LSST camera has 6 filters () and only five can be loaded in the filter exchange carousel. The current filter load strategy is to replace during daytime one of the or by the band at the start of dark time, when the lunar phase is lower than 30%. The process is reversed when the lunar phase is above 30%. We have performed a study using SNe Ia to assess the impact of the swap or on two metrics ( and N). The results are detailed in Appendix A . The conclusion is that the minimal loss is observed for the swap. This can be explained by the fact that the number of visits in the band is lower than the number of visits in the band. The decrease of the light curve signal-to-noise ratio due to the swap is lower if and bands are exchanged. Following the conclusion of this study we have swapped the and bands when the lunar phase is low in the design of the proposed strategies. The total number of () visits is distributed among the nights with a Moon phase lower (higher) than 30%.
@@ 5.3 WL metric
Season 1 results (Tab. 13) show that WL requirements are met for all fields except for ECDFS-a and ECDFS-b in the southern region. The number of visits in ,,, for season 1 is driven by WL requirements (Tab. 2). is thus close to one for these bands and for all fields but EDFS-a and EDFS-b. For these two fields the number of visits is a factor of 2 lower compared to the other DDFs (SCOC phase 2 recommendations) and the ratio is close to 0.5. For this season the requirement 16 (§2.2.2) is fulfilled for all the fields but EDFS-a and EDFS-b and only for and bands.
@@ 5.4 SNe Ia metric
We assume that the SNe Ia used to estimate SMoM have host galaxy redshifts measured from spectroscopic observations. We (conservatively) consider two spectroscopic resources contemporaneous with LSST to provide vital spectra and host redshifts, PFS/Subaru (Tamura et al., 2016) and 4MOST/TiDES (de Jong et al., 2019; Swann et al., 2019). The PFS/Subaru spectroscopic follow-up survey is designed to observe two of the northernmost LSST DDFs, COSMOS and XMM-LSS. About 20,000 host galaxy redshifts up to 0.8 will be measured after ten years. The 4MOST/TiDES survey is dedicated to the spectroscopic follow-up of extragalactic optical transients. About 50,000 host galaxy redshifts up to 1 will be collected for 5 years. This corresponds both to the WFD and the DD fields. The fraction of SNe Ia expected to have secure redshift measurements is taken from Mandelbaum et al. (2019) (Figure 1, right plot).
@@
Measuring cosmological parameters () with high accuracy (i.e. maximizing SMoM) requires to observe a large sample of well-measured SNe Ia at higher redshifts (Gris et al., 2023). This is a necessary but not sufficient condition to measure (,) with high accuracy because low-z SNe Ia are also required. Scenarios with UDFs lead (by design) to samples with 1000, 1500, 1800 SNe Ia with for DDF_DESC_0.80*, DDF_DESC_0.75*, DDF_DESC_0.70* observing strategies, respectively (Figure 12). Furthermore, a large fraction of SNe Ia in these samples (65-80, 40-60, 20-40, respectively) lead to accurate distance measurements ().